# Extremely Amenable Automorphism Groups of Countable Structures

### v1

@@ 5.2. Isomorphism types of $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$
6. Some Descriptive Set Theoretic Results
Since $\mbox{\rm Aut}(\mathbb{U}^{<}_{\Delta})$ are Polish groups that are (isomorphic to) closed subgroups of $S_{\infty}$, our results in this paper fit very well in the following research program promoted by Kechris–Nies–Tent [14].
Program.
  1. (1)
  2. (2)
Here $\mathcal{F}(S_{\infty})$ is the standard Borel space consisting of all closed subsets of $S_{\infty}$. The $\sigma$-algebra of Borel sets on $\mathcal{F}(S_{\infty})$ is generated (via iterations of countable union, countable intersection, and complementation) by sets of the form
$\{F\in\mathcal{F}(S_{\infty})\colon F\cap U\neq\emptyset\},$
where $U\subseteq S_{\infty}$ is a basic open set, i.e., of the form
$N_{\bar{a},\bar{b}}=\{g\in S_{\infty}\colon g(\bar{a})=\bar{b}\}$
for some tuples $\bar{a},\bar{b}$ of natural numbers of the same length. For other undefined terms and concepts, see [7].
We now focus on the class of all closed subgroups of $S_{\infty}$ that are extremely amenable. We denote this class by $\mathsf{EA}$. The following theorem shows that $\mathsf{EA}$ is a Borel subset of $\mathcal{F}(S_{\infty})$.
Theorem 6.1.
The class $\mathsf{EA}$ of extremely amenable closed subgroups of $S_{\infty}$ is Borel.
Proof.
$g\cdot(a_{1},\dots,a_{n})=(g(a_{1}),\dots,g(a_{n})).$
We note that the collection of all $\mathcal{L}$-structures with the universe $\mathbb{N}$ forms a Polish space $X_{\mathcal{L}}=\prod_{n,i\geq 1}X_{n,i}$, where each $X_{n,i}$ is a copy of the Cantor space $2^{\mathbb{N}^{n}}$. The map $G\mapsto\mathcal{A}_{G}$ from $\mathcal{F}(S_{\infty})$ to $X_{\mathcal{L}}$ is Borel. In fact, a subbasic open set in $X_{\mathcal{L}}$ is determined by a tuple $\bar{a}=(a_{1},\dots,a_{n})$ being assigned value 00 or value $1$. If $U$ is a basic open set in $X_{\mathcal{L}}$, then whether $\mathcal{A}_{G}\in U$ is determined by finitely many requirements of form $\exists g\in G\ g(\bar{a})=\bar{b}$ and its negation, where $\bar{a},\bar{b}$ are tuples of the same length. Thus $\{G\in\mathcal{F}(S_{\infty})\colon\mathcal{A}_{G}\in U\}$ is clearly a Borel set.
Let $\mathcal{K}$ be the age of $\mathcal{A}_{G}$. Then $\mathcal{A}_{G}=\mbox{Flim}\,\mathcal{K}$. By Theorem 2.2, $G=\mbox{\rm Aut}(\mathcal{A}_{G})=\mbox{\rm Aut}(\mbox{Flim}\,\mathcal{K})$ is extremely amenable iff $\mathcal{K}$ has the Ramsey property and every element of $\mathcal{K}$ is rigid. Thus to prove $\mathsf{EA}$ is Borel, it suffices to show both the following:
  1. (i)
  2. (ii)
Thus we have accomplished the first goal of the above-mentioned research program for the class of extremely amenable closed subgroups of $S_{\infty}$. Next we address the second goal and state some consequences of our results from the previous sections. We let $\cong_{\mathsf{EA}}$ denote the topological group isomorphism relation among groups in $\mathsf{EA}$. As noted in [14], it follows from Theorem 6.1 that $\cong_{\mathsf{EA}}$ is an analytic equivalence relation. This gives an upper bound in the Borel reducibility hierarchy for $\cong_{\mathsf{EA}}$. In the rest of this section we focus on lowers bounds for $\cong_{\mathsf{EA}}$.
@@ 6. Some Descriptive Set Theoretic Results
    $d_{i}\geq 0$ for all $i$; the set $\{i\in\mathbb{N}\colon d_{i}=0\}$ is infinite;
@@
Let $\mathsf{CDVS}_{0}$ consist of all sequences $(d_{i})_{i\in\mathbb{N}}\in\mathsf{CDVS}$ such that
$\inf\{d_{i}>0\colon i\in\mathbb{N}\}=0.$
Intuitively, $\mathsf{CDVS}$ consists of codes of all countable distance value sets, which are enumerated nonrepeatedly as sequences of real numbers. We include zero entries in the codes to accommodate finite distance value sets. Similarly, $\mathsf{CDVS}_{0}$ consists of codes of all countable distance value sets $\Delta$ with $\inf\Delta=0$. Both $\mathsf{CDVS}$ and $\mathsf{CDVS}_{0}$ are Borel subsets of $\mathbb{R}^{\mathbb{N}}$, and hence they are both standard Borel spaces. Recall the following definition of Borel reducibility among equivalence relations on standard Borel spaces.
@@
Let $E,F$ be equivalence relations on standard Borel spaces $X,Y$, respectively. We say that $E$ is Borel reducible to $F$, and denote it as $E\leq_{B}F$, if there is a Borel function $f\colon X\to Y$ such that for all $x,y\in X$,
$xEy\iff f(x)Ff(y).$
We say that $E$ and $F$ are Borel bireducible if both $E\leq_{B}F$ and $F\leq_{B}E$. We say that $E$ is strictly Borel reducible to $F$, and write $E<_{B}F$, if $E\leq_{B}F$ and $F\not\leq_{B}E$.
Next we define the equivalence relations of interest which are lower bounds for $\cong_{\mathsf{EA}}$ in the sense of $\leq_{B}$.

### v2

@@ 4.1. Isomorphism types of $\mbox{\rm Iso}(\mathbb{U}_{\Delta})$
5. The Complexity of the Equivalence between Distance Value Sets
In this section we study the complexity of the equivalence (see Definition 4.1) between countable distance value sets from the point of view of descriptive set theory. We first establish some notation.
@@ 5. The Complexity of the Equivalence between Distance Value Sets
    $d_{i}\geq 0$ for all $i$; the set $\{i\in\mathbb{N}\colon d_{i}=0\}$ is infinite; $d_{i}>0$ for some $i$;
@@
Intuitively, $\mathsf{CDVS}$ consists of codes of all countable distance value sets, which are enumerated as sequences of real numbers without repetitions. We include zero entries in the codes to accommodate finite distance value sets. $\mathsf{CDVS}$ is a Borel subset of $\mathbb{R}^{\mathbb{N}}$, and hence it is a standard Borel space. We also use the following notation:
$\begin{array}[]{lcl}\mathsf{CDVS}_{0}&=&\left\{(d_{i})_{i\in\mathbb{N}}\in\mathsf{CDVS}\colon\inf\{d_{i}>0\colon i\in\mathbb{N}\}=0\right\},\\ \mathsf{CDVS}_{+}&=&\left\{(d_{i})_{i\in\mathbb{N}}\in\mathsf{CDVS}\setminus\mathsf{CDVS}_{0}\colon\sup_{i}d_{i}<\infty\right\},\\ \mathsf{CDVS}_{\infty}&=&\left\{(d_{i})_{i\in\mathbb{N}}\in\mathsf{CDVS}\setminus\mathsf{CDVS}_{0}\colon\sup_{i}d_{i}=\infty\right\}.\end{array}$
These sets form a partition of $\mathsf{CDVS}$ into Borel subsets, and in particular they are all standard Borel spaces. $\mathsf{CDVS}_{0}$ consists of codes of all countable distance value sets $\Delta$ with $\inf\Delta=0$. Similarly, $\mathsf{CDVS}_{+}$ codes countable distance value sets $\Delta$ with $\inf\Delta>0$ and $\sup\Delta<\infty$, and $\mathsf{CDVS}_{\infty}$ codes those $\Delta$ with $\inf\Delta>0$ and $\sup\Delta=\infty$. In our discussions below, we will not distinguish a countable distance value set $\Delta$ from a code of $\Delta$ as an element of $\mathsf{CDVS}$.
@@
For any $A\subseteq\mathbb{R}^{\mathbb{N}}$, let $\approx_{A}$ denote $\approx\upharpoonright(A\times A)$.
Thus $\approx_{\mathsf{CDVS}}$ is the equivalence between countable distance value sets. Note that $\mathsf{CDVS}_{0}$, $\mathsf{CDVS}_{+}$ and $\mathsf{CDVS}_{\infty}$ are all invariant under the equivalence relation $\approx$.
In the rest of this section we will determine the exact complexity of $\approx_{\mathsf{CDVS}_{+}}$, $\approx_{\mathsf{CDVS}_{\infty}}$ and $\approx_{\mathsf{CDVS}}$ and give some upper and lower bounds for $\approx_{\mathsf{CDVS}_{0}}$ in the Borel reducibility hierarchy.
5.1. The exact complexity of $\approx_{\mathsf{CDVS}_{+}}$ and $\approx_{\mathsf{CDVS}}$
Lemma 5.1.
$\approx_{\mathsf{CDVS}}$ is Borel reducible to an isomorphism relation of countable structures. Consequently, $\approx_{\mathsf{CDVS}}$ is Borel reducible to $E^{\infty}_{S_{\infty}}$.
Proof.
$R^{\mathcal{A}}(d_{1},d_{2},d_{3})\iff\mbox{$(d_{1},d_{2},d_{3})$ is a $\Delta$-triangle.}$
It is clear that for $\Delta,\Lambda\in\mathsf{CDVS}$, we have $\Delta\approx\Lambda$ if and only if $\mathcal{A}_{\Delta}\cong\mathcal{A}_{\Lambda}$. This gives a Borel reduction from $\approx_{\mathsf{CDVS}}$ to an isomorphism relation of countable structures. ∎
Theorem 5.2.
The isomorphism relation of all countable linear orders is Borel reducible to $\approx_{\mathsf{CDVS}_{+}}$. Consequently, $E^{\infty}_{S_{\infty}}$ is Borel reducible to $\approx_{\mathsf{CDVS}_{+}}$.
Proof.
For a countable linear order $\mathcal{O}$, let $\psi(\mathcal{O})\subseteq(3,4)$ be an isomorphic copy of $\mathcal{O}$, and let
$\varphi(\mathcal{O})=\{2,4,6\}\cup\psi(\mathcal{O})\cup(2+\psi(\mathcal{O})).$
Then $\varphi(\mathcal{O})\in\mathsf{CDVS}_{+}$. We verify that $\varphi$ is a Borel reduction from the isomorphism relation of all countable linear orders to $\approx_{\mathsf{CDVS}_{+}}$.
First suppose $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are isomorphic countable linear orders. Then there is an order-preserving bijection $f$ between $\psi(\mathcal{O}_{1})$ and $\psi(\mathcal{O}_{2})$. For $x\in\varphi(\mathcal{O}_{1})$, let
$F(x)=\left\{\begin{array}[]{ll}f(x),&\mbox{ if $x\in\psi(\mathcal{O}_{1})$,}\\ f(x-2)+2,&\mbox{ if $x\in 2+\psi(\mathcal{O}_{1})$,}\\ x,&\mbox{ otherwise.}\end{array}\right$.
It is straightforward to check that $F$ is a bijection between $\varphi(\mathcal{O}_{1})$ and $\varphi(\mathcal{O}_{2})$ witnessing their equivalence. Thus $\varphi(\mathcal{O}_{1})\approx\varphi(\mathcal{O}_{2})$.
Conversely, suppose $\varphi(\mathcal{O}_{1})\approx\varphi(\mathcal{O}_{2})$ and this equivalence is witnessed by a bijection $g$. Let $\Delta=\varphi(\mathcal{O}_{1})$ and $\Lambda=\varphi(\mathcal{O}_{2})$, and let $\mathcal{A}_{\Delta}=(\Delta,R^{\Delta})$ and $\mathcal{A}_{\Lambda}=(\Lambda,R^{\Lambda})$ be the countable structures defined in the proof of Lemma 5.1. Then $g$ is an isomorphism witnessing $\mathcal{A}_{\Delta}\cong\mathcal{A}_{\Lambda}$.
Now note that $2$ is definable in both $\mathcal{A}_{\Delta}$ and $\mathcal{A}_{\Lambda}$ by the formula $\exists y\neg R(x,x,y)$. Thus $g(2)=2$. Moreover, $\psi(\mathcal{O}_{1})$ is definable in $\mathcal{A}_{\Delta}$ by the formula
$x\neq 2\wedge R(2,2,x)\wedge\exists y\neg R(2,x,y).$
Similarly, $\psi(\mathcal{O}_{2})$ is definable in $\mathcal{A}_{\Lambda}$ by the same formula. Thus $g$ sends $\psi(\mathcal{O}_{1})$ to $\psi(\mathcal{O}_{2})$. Finally, on $\psi(\mathcal{O}_{1})$, the $x<y$ relation is definable by the formula
$\exists z\,[\neg R(2,x,z)\wedge R(2,y,z)].$
Similarly for $\psi(\mathcal{O}_{2})$. Hence $g$ preserves the order << as a bijection between $\psi(\mathcal{O}_{1})$ and $\psi(\mathcal{O}_{2})$. It follows that the linear orders $\mathcal{O}_{1}$ and $\mathcal{O}_{2}$ are isomorphic.
It is routine to check that $\varphi$ is Borel. ∎
Corollary 5.3.
Both $\approx_{\mathsf{CDVS}_{+}}$ and $\approx_{\mathsf{CDVS}}$ are Borel bireducible with $E^{\infty}_{S_{\infty}}$. In particular, they are analytic and non-Borel.
Proof.
This follows immediately from Lemma 5.1 and Theorem 5.2, in view of the trivial fact that $\approx_{\mathsf{CDVS}_{+}}$ is Borel reducible to $\approx_{\mathsf{CDVS}}$ via the identity function. ∎
5.2. The exact complexity of $\approx_{\mathsf{CDVS}_{\infty}}$
Theorem 5.4.
Let $\Delta$ and $\Lambda$ be countable distance value sets with $\inf\Delta=\inf\Lambda=1$ and $\sup\Delta=\sup\Lambda=\infty$. Then $\Delta$ and $\Lambda$ are equivalent if and only if $\Delta=\Lambda$.
Proof.
The implication $\Leftarrow$ is obvious. We only show the $\Rightarrow$ direction. For this, let $f$ be a bijection between $\Delta$ and $\Lambda$ witnessing their equivalence.
For a subset $I\subseteq\mathbb{R}$, let $\Delta_{I}=\Delta\cap I$ and $\Lambda_{I}=\Lambda\cap I$. Let $\mathcal{A}_{\Delta}=(\Delta,R^{\Delta})$ and $\mathcal{A}_{\Lambda}=(\Lambda,R^{\Lambda})$ be the countable structures defined in the proof of Lemma 5.1. Then $f$ is an isomorphism witnessing $\mathcal{A}_{\Delta}\cong\mathcal{A}_{\Lambda}$.
First note that $\inf\Delta=1\in\Delta$ if and only if $\mathcal{A}_{\Delta}$ satisfies the sentence
$\exists x\,\forall y\,[R(x,x,y)\rightarrow\forall z\,R(z,z,y)].$
Similarly for $\Lambda$. Thus $1\in\Delta$ if and only if $1\in\Lambda$. We consider two cases.
Case 1: $1\in\Delta\cap\Lambda$. Since $\Delta$ and $\Lambda$ are distance value sets with $\sup\Delta=\inf\Delta=\infty$, we have that $k\in\Delta\cap\Lambda$ for all positive integers $k$. Note that $1$ is definable in $\mathcal{A}_{\Delta}$ by the formula
$\chi_{1}(x)=\forall y\,[R(x,x,y)\rightarrow\forall z\,R(z,z,y)].$
Similarly for $\mathcal{A}_{\Lambda}$. Thus $f(1)=1$. Also, we have that $\Delta_{[1,2]}$ is definable in $\mathcal{A}_{\Delta}$ by the formula
$\chi_{[1,2]}(x)=\forall y\,R(y,y,x).$
Similarly for $\Lambda_{[1,2]}$. Hence $f$ sends $\Delta_{[1,2]}$ to $\Lambda_{[1,2]}$.
Now for $k\geq 2$, define inductively
$\chi_{[1,k+1]}(x)=\exists y\,[\chi_{[1,k]}(y)\wedge R(1,y,x)].$
Then $\Delta_{[1,k+1]}$ is definable in $\mathcal{A}_{\Delta}$ by $\chi_{[1,k+1]}$, and similarly for $\Lambda_{[1,k+1]}$. It follows that $f$ sends $\Delta_{[1,k+1]}$ to $\Lambda_{[1,k+1]}$ for all $k\geq 1$.
Next we define, by induction on $k\in\mathbb{N}^{+}$ and $m\in\mathbb{N}$, a formula $\chi_{[1,1+\frac{k}{2^{m}}]}$ as follows.
$\begin{array}[]{rcl}\chi_{[1,1+\frac{k}{2^{0}}]}(x)&=&\chi_{[1,k+1]}(x)\mbox{ for all $k\geq 1$,}\\ \chi_{[1,1+\frac{k}{2^{m+1}}]}(x)&=&\forall y\,[R(x,x,y)\rightarrow\chi_{[1,1+\frac{k+2^{m}}{2^{m}}]}(y)].\end{array}$
For any $k\in\mathbb{N}^{+}$ and $m\in\mathbb{N}$, the formula $\chi_{[1,1+\frac{k}{2^{m}}]}$ defines $\Delta_{[1,1+\frac{k}{2^{m}}]}$ in $\mathcal{A}_{\Delta}$ and $\Lambda_{[1,1+\frac{k}{2^{m}}]}$ in $\mathcal{A}_{\Lambda}$. Hence $f$ sends $\Delta_{[1,1+\frac{k}{2^{m}}]}$ to $\Lambda_{[1,1+\frac{k}{2^{m}}]}$. This implies that for any dyadic rationals $1<a<b$, $f$ sends $\Delta_{[a,b]}$ to $\Lambda_{[a,b]}$. Hence $f$ must be the identity function, and so $\Delta=\Lambda$.
Case 2: $1\not\in\Delta\cup\Lambda$. The formula $\chi_{[1,2]}$ now defines $\Delta_{(1,2]}$ in $\mathcal{A}_{\Delta}$ and $\Lambda_{(1,2]}$ in $\mathcal{A}_{\Lambda}$. Thus $f$ sends $\Delta_{(1,2]}$ to $\Lambda_{(1,2]}$.
Let
$\varphi_{2}(x)=\exists y\,[\chi_{[1,2]}(y)\wedge\forall z\,(\chi_{[1,2]}(z)\rightarrow R(x,y,z))]\wedge\neg\chi_{[1,2]}(x)$
and let $D_{2}$ be defined by $\varphi_{2}$ in $\mathcal{A}_{\Delta}$. Then $D_{2}\subseteq\Delta_{(2,3]}$. We claim that $D_{2}\neq\varnothing$. To see this, let $a,b\in\Delta_{(1,2]}$ be such that $a<b$ and 2(a1)<b12(a-1)<b-1. Such $a,b$ exist since $\inf\Delta=1\not\in\Delta$. Then $2a\in\Delta$ since $\Delta$ is a distance value set, and $\varphi_{2}[2a]$ holds in $\mathcal{A}_{\Delta}$ as witnessed by $b$. Thus $2a\in D_{2}$. A similar argument shows $\inf D_{2}=2$. It follows that $\Delta_{(2,3]}$ is definable in $\mathcal{A}_{\Delta}$ by the formula
$\chi_{(2,3]}(x)=\forall y\,\forall z\,[(\chi_{[1,2]}(y)\wedge\varphi_{2}(z))\rightarrow R(x,y,z)]\wedge\neg\chi_{[1,2]}(x)$
and $\Delta_{(1,3]}$ is definable in $\mathcal{A}_{\Delta}$ by the formula
$\chi_{(1,3]}(x)=\chi_{[1,2]}(x)\vee\chi_{(2,3]}(x).$
Similarly for $\mathcal{A}_{\Lambda}$. Therefore $f$ sends $\Delta_{(2,3]}$ to $\Lambda_{(2,3]}$ and sends $\Delta_{(1,3]}$ to $\Lambda_{(1,3]}$.
Now for $k\geq 2$, define inductively
$\begin{array}[]{rcl}\varphi_{k+1}(x)&=&\exists y\,[\varphi_{k}(y)\wedge\forall z\,(\chi_{[1,2]}(z)\rightarrow R(x,y,z))]\wedge\neg\chi_{(1,k+1]}(x),\\ \chi_{(k+1,k+2]}(x)&=&\forall y\,\forall z\,[(\chi_{[1,2]}(y)\wedge\varphi_{k+1}(z))\rightarrow R(x,y,z)]\wedge\neg\chi_{(1,k+1]}(x),\\ \chi_{(1,k+2]}(x)&=&\chi_{(1,k+1]}(x)\vee\chi_{(k+1,k+2]}(x).\end{array}$
Then for all $k\geq 1$, $\chi_{(1,k+2]}$ defines $\Delta_{(1,k+2]}$ in $\mathcal{A}_{\Delta}$ and $\Lambda_{(1,k+2]}$ in $\mathcal{A}_{\Lambda}$. Hence for all $k\geq 2$, $f$ sends $\Delta_{(1,k+1]}$ to $\Lambda_{(1,k+1]}$.
The rest of the proof is similar to Case 1. We conclude that $f$ must be the identity function, and so $\Delta=\Lambda$. ∎
Theorem 5.5.
$\approx_{\mathsf{CDVS}_{\infty}}$ is Borel bireducible with $^{+}$.
Proof.
To each $\Delta\in\mathsf{CDVS}_{\infty}$ we associate a code for
$\Delta^{*}=\displaystyle\frac{\Delta}{\inf\Delta}.$
Then $\Delta^{*}\in\mathsf{CDVS}_{\infty}$ and $\inf\Delta^{*}=1$. By Theorem 5.4, the map $\Delta\mapsto\Delta^{*}$ is a Borel function witnessing that $\approx_{\mathsf{CDVS}_{\infty}}\,\leq_{B}\ =^{+}$.
Conversely, given any nonempty countable subset $A$ of the interval $(0,1)$, let $\Delta_{A}$ be the subsemigroup of $(\mathbb{R}^{+},+)$ generated by $\{1\}\cup(1+A)$. Then $\Delta_{A}$ is a countable distance value set with $\inf\Delta_{A}=1$ and $\sup\Delta_{A}=\infty$. Thus $\Delta_{A}\in\mathsf{CDVS}_{\infty}$. Moreover, $\Delta_{A}\cap(1,2)=1+A$. By Theorem 5.4, if $A$ and $B$ are two nonempty countable subsets of $(0,1)$, then $\Delta_{A}$ and $\Delta_{B}$ are equivalent if and only if $\Delta_{A}=\Delta_{B}$, if and only if $1+A=1+B$, and if and only if $A=B$. Thus $A\mapsto\Delta_{A}$ is a Borel function witnessing that $^{+}\,\leq_{B}\ \approx_{\mathsf{CDVS}_{\infty}}$. ∎
5.3. The complexity of $\approx_{\mathsf{CDVS}_{0}}$
In this subsection we give some upper and lower bounds for the complexity of $\approx_{\mathsf{CDVS}_{0}}$. We also discuss some related equivalence relations.
Theorem 5.6.
Let $\Delta$ and $\Lambda$ be countable distance value sets with $\inf\Delta=\inf\Lambda=0$. Then the following are equivalent:
  1. (1)
    $\Delta$ and $\Lambda$ are equivalent.
  2. (2)
    There is a positive real number $r$ such that $\Lambda=r\cdot\Delta$.
Proof.
[…]
Fix a sequence $(\delta_{i})_{i\in\mathbb{N}}$ in $\Delta$ such that $\delta_{i+1}<{\delta_{i}}/{2}$ for all $i$. Such a sequence exists since $\mathcal{A}_{\Delta}$ satisfies the sentence $\forall x\,\exists y\,\neg R(y,y,x)$. Since $f$ is an isomorphism, we have $f(\delta_{i+1})<{f(\delta_{i})}/{2}$ for all $i$.
We claim that $f$ is continuous as a function from $\Delta\subseteq\mathbb{R}$ to $\Lambda\subseteq\mathbb{R}$. To see this, fix $x\in\Delta$. For any $\epsilon>0$, find an $i$ such that $\delta_{i}<x$ and $f(\delta_{i})<\epsilon$. For any $y\in\Delta$ with $|x-y|<\delta_{i}$, $(x,y,\delta_{i})$ is a $\Delta$-triangle. Thus $(f(x),f(y),f(\delta_{i}))$ is a $\Lambda$-triangle. In particular, $|f(x)-f(y)|<f(\delta_{i})<\epsilon$. The actually proves that for all positive $\delta$, $f$ is uniformly continuous on $\Delta_{>\delta}\triangleq\Delta\cap(\delta,+\infty)$.
Next we show that $f$ is in fact linear. In the proof we use notation such as $\Delta_{>\delta}$, $\Delta_{<\delta}$, and $\Delta_{[a,b]}$ with obvious meanings.
To show that $f$ is linear, assume $x,y,x+y\in\Delta$. We claim that
  1. (a)
    $f(\Delta_{[|x-y|,x+y]})\subseteq\Lambda_{[|f(x)-f(y)|,f(x)+f(y)]}$;
  2. (b)
[…]
  3. (c)
[…]
Let $z\in\Delta$. (a) is equivalent to saying that if $(x,y,z)$ is a $\Delta$-triangle, then $(f(x),f(y),f(z))$ is a $\Lambda$-triangle, which is immediate from our assumption. Also immediate is that if $z>x+y$, then $(x,y,z)$ is not a $\Delta$-triangle. Thus $(f(x),f(y),f(z))$ is not a $\Lambda$-triangle, and so either $f(z)>f(x)+f(y)$ or f(z)<|f(x)f(y)|f(z)<|f(x)-f(y)|. Similarly for z<|xy|z<|x-y|. Now, to prove (b), assume toward a contradiction that there are $z_{1},z_{2}\in\Delta$ with $z_{1},z_{2}>x+y$ such that $f(z_{1})<|f(x)-f(y)|$ and $f(z_{2})>f(x)+f(y)$. Without loss of generality assume $z_{1}<z_{2}$ (the argument for the case $z_{1}>z_{2}$ is similar). By the continuity of $f$ and the density of both $\Delta$ and $\Lambda$, we may find $z_{1}=a_{1},\dots,a_{n}=z_{2}$ in $\Delta$ such that $a_{k}>x+y$ and $|f(a_{k+1})-f(a_{k})|<f(x)+f(y)-|f(x)-f(y)|$ for all $1\leq k\leq n$. Then there must be an $a_{k}$ with $|f(x)-f(y)|\leq f(a_{k})\leq f(x)+f(y)$, contradicting $a_{k}>x+y$. This proves (b). A similar argument for $z_{1},z_{2}<|x-y|$ proves (c).
Since $f$ is a bijection from $\Delta$ to $\Lambda$, the claimed inclusions are actually equalities. By considering sufficiently small $\delta_{i}$, we see that $f(\Delta_{<|x-y|})=\Lambda_{<|f(x)-f(y)|}$, and thus $f(\Delta_{>x+y})=\Lambda_{>f(x)+f(y)}$. By the continuty of $f$ again, we have $f(x+y)=f(x)+f(y)$.
Finally, we argue that for any $x,t\in\Delta$, we have $f(x)/x=f(t)/t$. This proves (2).
For each $k>0$, find $0<\delta_{k}<1/2^{k}$ by the continuity of $f$ such that for all $z\in\Delta$, $|t-z|<\delta_{k}$ implies |f(t)f(z)|<1/k|f(t)-f(z)|<1/k. Now for each $k$, find positive rational number $n_{k}/m_{k}$ such that $0<t-(n_{k}/m_{k})x<\delta_{k}/2$. Note that by the density of $\Delta$ and bijectivity of $f$, $\delta_{k}\to 0$ as $k\to\infty$. Thus $n_{k}/m_{k}\to t/x$ as $k\to\infty$.
Letting $k\to\infty$, we get $f(t)=(t/x)f(x)$ and thus $f(x)/x=f(t)/t$. ∎
In view of Theorem 5.6, we define the following notion.